OpenStudy (anonymous):

write a slope-intercept equation of the line whose is parallel to y=4x-2 and has the y intercept of (0,6)

OpenStudy (luigi0210):

Since it's parallel we have to use the same slope from the equation which is 4. And we use the point (0,6) Just plug it into y-y1=m(x-x1) The 0 is the x1 The 6 is the y1 and the m is 4

OpenStudy (luigi0210):

So the equation would be y-6=4(x-0) y-6=4x y=4x-6

OpenStudy (anonymous):

Thank you you broke that down so I can even understand it

OpenStudy (luigi0210):

No prob